complete the square
9x=3x eponent 2 -1

Answer by stanbon(75887)

complete the square
9x=3x exponent 2 -1
---------
Rearrange:
3x^2 - 9x - 1
---
3(x^2 - 3x + ?) = 1 + 3*?
---
3(x^2 - 3x + (3/2)^2) = 1 + 3(3/2)^2
---
3(x-(3/2))^2 = 1 + 3(9/4)
3(x-(3/2))^2 = (31/4)
(x-(3/2))^2 = (31/12)
------------------------
If you want to solve this take the square root of both sides
==========================================
Cheers,
Stan H.